(TCO 3) Find the acceleration produced by a total force of 175 N on a mass of 5.40 kg.
Answer #1

Using the second law of Newton,

F = ma....here F is the total force, m is the mass and a is the acceleration

Or, a = F/m = 175 /5.40 = 32.407 m/s2